on the basis of the following production function Q=AK4/5L1/5 write down the general mathematical form of the demand function deriving from this form of technology in the case of a profit maximizing firm
1
Expert's answer
2018-05-13T07:18:08-0400
Q=AK4/5L1/5
A linear demand curve can be plotted using the following equation. Qd = a – b(P).
The inverse demand equation can also be written as P = a -b(Q), where a = intercept where price is 0, b = slope of demand curve.
So, P = a - b*AK4/5L1/5.